OpenStudy (anonymous):

Which is equal to 4^2 • 4^8?

OpenStudy (anonymous):

\[a. 4^16 \]

OpenStudy (igreen):

When multiplying exponents with the same base, keep the base and add the exponents. What do you get? @Jennaluv

OpenStudy (anonymous):

The 16 is part of the exponent.

OpenStudy (igreen):

\(4^2 \cdot 4^8 = 4^{2+8} = ~?\)

OpenStudy (anonymous):

\[b.4^10\]The 0 is part of the exponent.

OpenStudy (anonymous):

What iGreen says :) it's another property of of exponents 

OpenStudy (igreen):

\(4^2 \cdot 4^8 = 4^{2+8} = ~4^{10}\)

OpenStudy (anonymous):

c.\[c. 4^6\]

OpenStudy (anonymous):

d.\[4^-6 \] The 6 is part of the exponent. The exponent is -6... x3

OpenStudy (anonymous):

I am so sorry @iGreen , I didn't see your replies... So to multiply these all I have to do is add the exponents, and keep my number, and I have my answer?

OpenStudy (anonymous):

Yep! As long as they have the same base (b), then you can add them when multiplying the terms

OpenStudy (igreen):

Yup, that's it.

OpenStudy (anonymous):

Well, both of you now have 2 medals! yay.... cx

OpenStudy (anonymous):

Thanx guys!

OpenStudy (anonymous):

Could you help me with another?

OpenStudy (igreen):

Sure.

OpenStudy (anonymous):

Il post another thing so you can get another medal.